# EvoXplain: When Machine Learning Models Agree on Predictions but Disagree on Why Measuring Mechanistic Multiplicity Across Training Runs

### v1

@@ 8 Discussion
EvoXplain shows that even models with high prediction accuracy do not guarantee explanation stability, even when datasets are thought to be ”simple”. Across both datasets, multiple model classes resulted in multiple explanatory modes with considerable support. This suggests that interpretability cannot be inferred from model form or accuracy alone. For example, linear models commonly assumed to be transparent do not consistently converge to a single explanatory mechanism under standard retraining procedures. In our experiments, Logistic Regression exhibits explanation multiplicity across datasets when regularisation is allowed to vary, despite high and stable predictive accuracy (lipton2018myth; rudin2019stop). This challenges the common practice of treating transparency and accuracy as sufficient conditions for interpretability. The goal is not to eliminate variability, but to characterise the range of mechanisms that arise under standard retraining procedures commonly used in practice.
A dominant operational assumption in applied machine learning is that inspecting a single trained model instance is sufficient to characterise a system’s reasoning. EvoXplain clarifies the limits of that assumption, as it shows that models with the same accuracy can have different internal reasoning structures. This implies that one attribution profile is likely just one out of many possible ones. Accordingly, interpretability should be framed as a population-level property of trained models rather than as an attribute of a single fitted instance. While predictive performance remains stable, explanation multiplicity matters because explanations are often used to justify, audit, or regulate model-driven decisions at the individual level. This suggests that, rather than just focusing on a single model instance, we need to develop stability-aware interpretability of models that provides explanations based on the behaviour of models over multiple instances or runs (arrieta2020explainable; selbst2018intuitive).
@@
Therefore, inconsistency between explanations is not something to be averaged out as noise. Instead, we reframe it as an empirical observation. In this sense, explanation multiplicity is a property to be measured, reported, and reasoned about, rather than suppressed. In the presence of multiple modes of explanation, the decision to select one becomes a question of science or governance, rather than something the training pipeline can resolve automatically (ecegai2023aiact; rudin2019stop).
@@
This is particularly important because explanations have immediate consequences in certain domains. In criminal justice contexts, multiple basins of attraction counteract the assumption that a model provides a singular, stable justification umbach2023predictive. In biomedicine, instability observed even on standard datasets warns that attribution-based interpretations should be treated with caution when used to support clinical claims agrawal2023breast. Thus, EvoXplain helps to identify cases where a single explanation is insufficient to account for the behaviour of a model under retraining. In both settings, explanation stability becomes relevant precisely because explanations are used as part of decision-making processes rather than as post-hoc descriptions.
@@
In our experiments, we focus on tabular learning, where repeated training is manageable. However, retraining large language models (LLMs) presents a more difficult challenge. We do not claim that the specific geometric structures observed here generalise directly to large-scale models; rather, we argue that the underlying issue of mechanistic non-identifiability is likely to be amplified in such settings. LLMs train in very high-dimensional, non-convex spaces, are exposed to stochastic data, and follow fine-tuning paths that may diverge substantially. Whether models that exhibit similar external behaviour in such settings rely on the same internal mechanisms remains an open empirical question. Measuring this at scale is challenging, but not impossible: models and their fine-tuned variants can be compared, checkpoints sampled, and attribution patterns analysed across independent runs (olah2020circuits). EvoXplain provides a conceptual starting point. The instability observed even in small models suggests that non-identifiability in foundation models may be important and deserves further study (lakshminarayanan2017simple).
@@
There are boundaries to this work. We analyse explanation stability using attribution methods; other interpretability approaches, such as causal abstraction or circuit tracing, may offer different perspectives (olah2020circuits). EvoXplain does not attempt to correct bias, determine which basin is “correct”, or quantify downstream risks such as fairness or failure modes selbst2018intuitive. We consider only a limited number of datasets and model classes; this study represents an initial step toward broader analysis across architectures, attribution schemes, and domains such as reinforcement learning or language modelling. These limitations define the scope of the present study rather than its validity. The contribution of EvoXplain is not to propose a new explainer, but to provide a diagnostic framework that makes explanation instability measurable, comparable, and reportable across model populations.
Despite these limitations, EvoXplain opens new research questions. Are explanation basins correlated with latent subgroups, causal mechanisms, or inductive bias pathways (marx2020predictive)? Is it possible to prefer certain basins for safety, fairness, or generalisability (rudin2019stop)? More generally, EvoXplain shifts interpretability from analysing individual models to analysing populations of models trained under the same procedure. This reframing allows explanation stability to be measured, compared, and monitored at the level of model populations.

### v2

@@ 8 Discussion
EvoXplain shows that even models with high prediction accuracy do not guarantee explanation stability, even when datasets are thought to be ”simple”. Across both datasets, multiple model classes resulted in multiple explanatory modes with considerable support. This suggests that interpretability cannot be inferred from model form or accuracy alone. For example, linear models commonly assumed to be transparent do not consistently converge to a single explanatory mechanism under standard retraining procedures. In our experiments, Logistic Regression exhibits explanation multiplicity across datasets when regularisation is allowed to vary, despite high and stable predictive accuracy (Lipton (2018); Rudin (2019)). This challenges the common practice of treating transparency and accuracy as sufficient conditions for interpretability. The goal is not to eliminate variability, but to characterise the range of mechanisms that arise under standard retraining procedures commonly used in practice.
A dominant operational assumption in applied machine learning is that inspecting a single trained model instance is sufficient to characterise a system’s reasoning. EvoXplain clarifies the limits of that assumption, as it shows that models with the same accuracy can have different internal reasoning structures. This implies that one attribution profile is likely just one out of many possible ones. Accordingly, interpretability should be framed as a population-level property of trained models rather than as an attribute of a single fitted instance. While predictive performance remains stable, explanation multiplicity matters because explanations are often used to justify, audit, or regulate model-driven decisions at the individual level. This suggests that, rather than just focusing on a single model instance, we need to develop stability-aware interpretability of models that provides explanations based on the behaviour of models over multiple instances or runs (Arrieta and others (2020); Selbst and Barocas (2018)).
@@
Therefore, inconsistency between explanations is not something to be averaged out as noise. Instead, we reframe it as an empirical observation. In this sense, explanation multiplicity is a property to be measured, reported, and reasoned about, rather than suppressed. In the presence of multiple modes of explanation, the decision to select one becomes a question of science or governance, rather than something the training pipeline can resolve automatically (European Commission (2023); Rudin (2019)).
This perspective aligns with recent safety assurance methodologies for machine learning, such as AMLAS and the “BIG Argument” for AI safety proposed by Habli and colleagues (Hawkins et al. (2021); Habli et al. (2025)), which emphasise the need for explicit evidence of model robustness beyond predictive accuracy alone. Within such frameworks, explanation stability is not merely an interpretability preference but a form of verification evidence. EvoXplain contributes precisely this type of evidence by quantifying how model reasoning varies under standard retraining procedures, thereby supporting or challenging claims of reliability in safety-critical ML systems (Paterson et al. (2025)).
A closely related operational response to model multiplicity is to aggregate explanations over a set of near-optimal models and report uncertainty around the aggregate. For instance, Rashomon-set approaches in AutoML construct a set of models within a performance tolerance and then average partial dependence profiles across that set to obtain a single ”Rashomon PDP”, optionally accompanied by confidence intervals to visualise disagreement Cavus et al. (2025). EvoXplain highlights a stronger failure mode: when explanation variability forms separable basins of near-equivalent accuracy, aggregation is not merely lossy but can be methodologically misleading, producing a ghost explanation that does not correspond to any model in the population. In such settings, interpretability requires representing the set of mechanistic regimes and their relative support (frequency), rather than collapsing them into a single averaged explanation with uncertainty bands.
@@
This is particularly important because explanations have immediate consequences in certain domains. In criminal justice contexts, multiple basins of attraction counteract the assumption that a model provides a singular, stable justification Umbach et al. (2023). In biomedicine, instability observed even on standard datasets warns that attribution-based interpretations should be treated with caution when used to support clinical claims Agrawal and Jain (2023). Thus, EvoXplain helps to identify cases where a single explanation is insufficient to account for the behaviour of a model under retraining. In both settings, explanation stability becomes relevant precisely because explanations are used as part of decision-making processes rather than as post-hoc descriptions.
@@
In our experiments, we focus on tabular learning, where repeated training is manageable. However, retraining large language models (LLMs) presents a more difficult challenge. We do not claim that the specific geometric structures observed here generalise directly to large-scale models; rather, we argue that the underlying issue of mechanistic non-identifiability is likely to be amplified in such settings. LLMs train in very high-dimensional, non-convex spaces, are exposed to stochastic data, and follow fine-tuning paths that may diverge substantially. Whether models that exhibit similar external behaviour in such settings rely on the same internal mechanisms remains an open empirical question. Measuring this at scale is challenging, but not impossible: models and their fine-tuned variants can be compared, checkpoints sampled, and attribution patterns analysed across independent runs (Olah et al. (2020)). EvoXplain provides a conceptual starting point. The instability observed even in small models suggests that non-identifiability in foundation models may be important and deserves further study (Lakshminarayanan et al. (2017)).
@@
There are boundaries to this work. We analyse explanation stability using attribution methods; other interpretability approaches, such as causal abstraction or circuit tracing, may offer different perspectives (Olah et al. (2020)). EvoXplain does not attempt to correct bias, determine which basin is “correct”, or quantify downstream risks such as fairness or failure modes Selbst and Barocas (2018). We consider only a limited number of datasets and model classes; this study represents an initial step toward broader analysis across architectures, attribution schemes, and domains such as reinforcement learning or language modelling. These limitations define the scope of the present study rather than its validity. The contribution of EvoXplain is not to propose a new explainer, but to provide a diagnostic framework that makes explanation instability measurable, comparable, and reportable across model populations.
Despite these limitations, EvoXplain opens new research questions. Are explanation basins correlated with latent subgroups, causal mechanisms, or inductive bias pathways (Marx et al. (2020))? Is it possible to prefer certain basins for safety, fairness, or generalisability (Rudin (2019))? More generally, EvoXplain shifts interpretability from analysing individual models to analysing populations of models trained under the same procedure. This reframing allows explanation stability to be measured, compared, and monitored at the level of model populations.